how should we calculate molecular formula when vapour density does not given?

Grade:12

1 Answers

Ramesh V
14 years ago

To find the no. of moles use the formula:

n=m/MM or moles= mass divided by Molecular mass

At STP conditions , we have many relations ,

     If Volume is known at stp for gas.
     Mole = Volume / 22.4 liters

     If no. of molecules/ions/atoms... are known
     Mole = no. of substance / (6.023 * 10^23)

     If mass , and formula mass are known
     Mole = Mass / Formula Mass

     If vapour density and mass are known
     Mole = Mass / (Vapour Density * 2)            etc.

If u r question is to calculate no. of moles if vapour density is given, then

molecular formula : A chemical formula that shows the number and kinds of atoms in a molecule.

MM = dRT/P
I would like to explain through an example:
Q) A certain compound containing only carbon and hydrogen was found to have a vapor density of 2.550 g/L at 100°C and 760 mm Hg. If the    empirical formula of this compound is CH, what is the molecular formula of this compound?

Ans) use the formula: MM = dRT/P
100°C = 373K
760 torr = 1atm
MM = (2.550g/L)(0.08206/mol*K)(373K)/(1atm)
MM = 78.05g/mol
MM CH = 13 g/mol
78.05g/mol / 13g/mol = 6

compund is :C6H6

--

Please feel free to post as many doubts on our disucssion forum as you can. If you find any question difficult to understand - post it here and
we will get you the answer and detailed solution very quickly.We are all IITians and here to help you in your IIT JEE preparation. All the best.

Regards,
Naga Ramesh
IIT Kgp - 2005 batch